A simple random sample of size n drivers were asked if they drive a car manufactured in a certain country. Of the drivers​ surveyed, responded that they did. Determine if more than half of all drivers drive a car made in this country at the level of significance. Complete parts ​(a) through ​(c).
​(a) Determine the null and alternative hypotheses.
​:
p
equals

​:
p
greater than
​(b) State the conclusion for the test.
Choose the correct answer below.
A.
because the​ P-value is the level of significance.
B.
because the​ P-value is the level of significance.
C.
because the​ P-value is the level of significance.
D.
because the​ P-value is the level of significance.
​(c) State the conclusion in context of the problem.
1
Expert's answer
2020-12-20T18:38:48-0500

a) The following null and alternative hypotheses need to be tested:


"H_0: p\leq0.5"

"H_1: p>0.5"

b) The z-statistic is computed as follows:


"z=\dfrac{\bar{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

Let "n=200, x=115"


"\bar{p}=\dfrac{x}{n}=\dfrac{115}{200}=0.575"

Then


"z=\dfrac{\bar{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

"z=\dfrac{0.575-0.5}{\sqrt{\dfrac{0.5(1-0.5)}{200}}}\approx2.1213"

Using the P-value approach: The p-value is "p=P(Z>2.1213)=0.016948," and since "p=0.016948<0.05=\alpha," it is concluded that the null hypothesis is rejected.


c) There is enough evidence to claim that the population proportion "p" is greater than "p_0=0.5," at the "\alpha=0.05"  significance level.